Hello,

I need help solving this problem:

A rectangle is 6 centimeters wide and 9 centimeters long. If the length and width are increased by the same amount, the area is increased by 100 square centimeters. Find the new dimensions.

Thank you

Answer by scott8148(6628)

9*6=54 ... (9+x)(6+x)=54+100 ... 54+9x+6x+x^2=154 ... x^2+15x-100=0

factoring gives (x+20)(x-5)-0 ... x+20=0 ... x=-20 ... x-5=0 ... x=5

new dimensions are 6+5=11 and 9+5=14